OpenStudy (anonymous):

The volume of a square based pyramid is eighty-one cubic meters. What is the volume of the cube with the same side length and height?

OpenStudy (anonymous):

V = (1/3)b^2h where b = base length, h = height from vertex to the base V = 13 b^3 because base and height are the same Since the cube has volue of 81 m^3, then b^3 = 81 b = 81^(1/3) V = 13 ( 81^(1/3) ) V = (1/3)(81) V = 27 m^3

OpenStudy (anonymous):

do you understand how i got that? @tparker942014

OpenStudy (anonymous):

no

OpenStudy (anonymous):

lol okayy what part are you confused on?

OpenStudy (anonymous):

yeah i got it thank you

OpenStudy (anonymous):

i just had to read it over i got confused with the multiply signs

OpenStudy (anonymous):

ur welcome!! (: let me know if you need anymore help..

OpenStudy (anonymous):

and yea that happens!!

OpenStudy (anonymous):

i have one more

OpenStudy (anonymous):

A rectangular pizza is cut into 6 pieces along the length and 4 pieces along the width. If you place 3 tomato slices on each piece of pizza, how many tomato slices will you use?

OpenStudy (anonymous):

okayy!!

OpenStudy (anonymous):

is it 30?

OpenStudy (anonymous):

do you have answer choices?

OpenStudy (anonymous):

okay gimme a sec to figure this out...

OpenStudy (anonymous):

okay its really simple... you can do it one of two ways. ready?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

okay so you can either do it by 6*4*3 or 6*3 then *4 or reverse 6*4*3 and then what do you get?

OpenStudy (anonymous):

6*4 and then *3*******

OpenStudy (anonymous):

thanks thats really easy

OpenStudy (anonymous):

you're welcome!! see i told youu...... anymore? if not, just @ my name and ill be there!!

OpenStudy (anonymous):

yes i have another one

OpenStudy (anonymous):

go right ahead

OpenStudy (anonymous):

What is the volume of the composite figure?

OpenStudy (anonymous):

i honestly dont even know on this one?!?! lemme find you some help.

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@e.mccormick